OpenStudy (anonymous):

Find the three terms in the sequence after the last one given. (Also with directions on how to find it plz) 1. 1,4,16,64 A.512 2048 8192 B. 128 512 2048 C. 256 1024 4096 D. 1024, 4096, 16384

OpenStudy (dan815):

4^1 =?

OpenStudy (dan815):

4^2=?

OpenStudy (dan815):

4^3=?

OpenStudy (dan815):

1 4 16 64 *4 *4 *4 *4

OpenStudy (anonymous):

@dan815 so what i think your trying to say is that there is a pattern like that and just find & ill have my problem solved ?

OpenStudy (dan815):

yes

OpenStudy (dan815):

can u see how u are multiplying by 4 each number

OpenStudy (dan815):

to get the 2nd

OpenStudy (dan815):

1*4=4 4*4=16 16*4=64 64*4=?

OpenStudy (anonymous):

@dan815 yes i get it now, i didnt know how to solve it but now i know what to do i can do my worksheet ! Thanks :)

OpenStudy (dan815):

ok good

OpenStudy (dan815):

good job!!! work hard have fun@